OpenStudy (anonymous):

Part 1: Make up an angle measure for angle C and a length for one of the sides (AB, BC, or CA). Part 2: Use a trigonometric ratio (sine, cosine, or tangent) to solve for one of the other side lengths. For full credit, please provide both the ratio and work used to solve for the missing side length. Round your solution to the nearest hundredth.

OpenStudy (anonymous):

So confused :(

OpenStudy (anonymous):

Okay so, 1. We start by measuring the length of an angle with a protractor (if that's what you are allowed to use). Measure, for example, angle C. 2. Angle C should be 45 degrees. 3. Now we wish to measure the side lengths. Check the length of BC, should be 4.5 cm 4. AB = BC = tan45 (tangent of angle C which is 45 degrees) = 4.5 cm 5. AC =BC/cos(angle C) = 4.5/cos45 = 6.36cm 6. Take 4.5/Cos45 = approximately 6.36396 cm (round it to the hundredth, should be around 6.364) 7. For angle CAB, the measure is 45 degrees too because 180 - 90 - 45 = 45 by angle sum property of a triangle. And that's all you need. Credits to Callisto for his explanation; this problem has been solved earlier. A required figure wasn't attached here though. Regards, Ven

OpenStudy (anonymous):

THANK YOU SO MUCH :) im new so umm how do i give you the medal thing ? LOL